Greenhab Report – November 28th

Crew 216 GreenHab Report 28-NOV-19
Crew GreenHab Officer: Michael Ho

Environmental control: Door opened for ventilation at 12.05pm, closed at 5.30 pm.
Average temperature: 20C
Hours of supplemental light: Automatic system
Daily water usage for crops: 20L
Daily water usage for research and/or other purposes: nil
Water in Blue Tank – enough
Time(s) of watering for crops: 09.30 am; 4.40 pm.

(Established plants introduced since 03/11/2019 till 22/11/2019 please refer to Crew 215 GreenHab Report of 22/11/2019.)

Harvest: None
Support/supplies requested: None
The 3 white round trays of snow peas are now hanging nicely! Strawberry nearest the door seems to be in the poorest condition of the 3, while the one furthest from the door seems to be in the best. Am wondering whether the hot air from the heater is the cause. Am moving the long tray right in front of the door during door opening hours so that they enjoy the cool air during that time. Will monitor over the next few days.
Inventory of all pots of plantings done. Paper is on the table.

Commander Report – November 28th

Crew 216 Commander Report 28-NOV-2019
Sol: 4
Summary Title: A Martian Thanksgiving
Author’s name: Marc Levesque
Crew 216 spent a low key day after three busy days taking advantage of good weather to accomplish six EVAs in support of projects. The morning saw an EVA team cleaning the muddy rovers, while the afternoon was devoted to preparing and eating our holiday meal. Our head chef Andrew started by putting a chicken into the convection oven for its 2.5 hour roasting. He then went to work on making biscuit and two fruit pies. In the meantime, Marc cut up Watney potatoes to join the roasting chicken. Rich pitched in with making the stuffing to round out the menu. Evgenia filmed the entire process so that she could add a homey touch to her documentary film. All hands then pitched in for the massive clean up. After some time to digest the sumptuous meal and crank out the day’s reports, the crew will settle into it’s second movie night. Last night was the premier of The Martian (of course), with tonight’s movie to be Interstellar. Our crew are movie buffs, especially of the space genre.

Research Report – November 28th

Crew 216 Research Report 28-Nov-2019
Sol 4
Submitted by: Whittle

1) Psych Scale (TAMU IRB2019-1042D) – Nothing further to report.

2) Medical Casts – Continuation of Crew 215 research. Nothing further to report.

3) In-Situ Fluorescent Mineral Prospecting – Nothing further to report.

4) MDRS Mapping – Nothing further to report (awaiting return of map with desired annotations).

5) MDRS Communications – Nothing further to report.

Misc: Crew demonstrated ability to cook a passable thanksgiving meal in the Hab.

Operations Report – November 28th

Crew 216 Operations Report 28-Nov-19
SOL: 4
Name of person filing report: Mike Lawson
Non-nominal systems: RAM rover power charging stations inoperable, electrical issue, ongoing.
Notes on non-nominal systems:
Generator: run
Hours run: 12.3
From what time last night: 1902
To what time this morning: 0723
List any additional daytime hours when the generator was run: N/A Solar— SOC % (Before generator is run at night): 69 %
Diesel Reading – 40
Station Propane Reading – 58
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1462347
Water (static tank) (gal): 375
Static to Loft Pump used – yes
Water in Green Hab (gal): 267
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: no
Hours: 114.5
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 100
Currently charging:
Opportunity rover used: no
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes
Curiosity rover used: no
Hours: 119.6
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 100
Currently charging: yes
Notes on rovers: will have to juggle charging sequence.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where?
CrewCar used and why, where?
General notes and comments:
Summary of internet: No serious issues.
Summary of suits and radios: Power strip to radio charger farm was inadvertently switched off, setting low power alarm off on one radio. Reset, no further issues.
Summary of Hab operations: Routine maintenance and housekeeping. Summary of GreenHab operations: Routine watering.
Summary of Science Dome operations: None, crew holiday.
Summary of RAM operations: RAM Rover power charging stations inop, electrical issues remain unresolved. RAM used to fabricate two baking pans from residue food cans for use in preparing holiday meal. Rover maintenance EVA performed.
Summary of any observatory issues: None, crew holiday.
Summary of health and safety issues: None.
Questions, concerns and requests to Mission Support: 1 bag NB trash in rear airlock. Request that a knife sharpener be added to the kitchen set for future mission. Crew used a field rock for sharpening. expedient (obsidian sample) for immediate need.

EVA Report – November 28th

Crew 216 EVA Report 28-Nov-2019
EVA # 07
Author: Whittle
Purpose of EVA: Rover maintenance
Start time: 1015
End time: 1130
Narrative: Crew cleaned Opportunity, Spirit, and Curiosity. Whilst outside, crew secured outside panel on Hab wall by water tank which had been flapping and banging.
Destination: Hab
Coordinates (use UTM NAD27 CONUS): 12 S 0518227 4250721.
Participants: Whittle, Kennedy, Ho, Alexandrova
Road(s) and routes per MDRS Map:
1) In vicinity of Hab
Mode of travel: Foot

Sol Summary – November 28th

Crew 216 Sol Summary Report 28-NOV-19
Sol: 4
Summary Title: Holiday on Mars
Author’s name: Marc Levesque
Mission Status: Active
Sol Activity Summary: A morning EVA cleaned the rovers which had become quite muddy from several EVAs accomplished during the hectic first three days of activity. Following that, the crew focused on preparing a lavish-for-Mars sit down dinner that featured roasted chicken and potatoes, stuffing, biscuits, and two fruit pies, with the blueberry/apple taking top prize.

Look Ahead Plan: Anticipated wet weather will curtail any EVA plans, but it will provide an opportunity for filmed interviews of crew members.

Anomalies in work: N/A.
Weather: Cold and windy overnight.
Crew Physical Status: Healthy
EVA: One EVA conducted.
Reports to be filed: Sol Summary Report, Commander Report, Operations Report, GreenHab Report, Journalist Report, Research Report, EVA Report, and Photos.
Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site